Calculate entropy change if 2 kg of water at 331 K is heated to 333 K at 1 bar pressure. The specific heat capacity of water is 4200J/kg.K.
1
Expert's answer
2020-12-09T09:43:00-0500
"\Delta S=mc_p\ln{\frac{T_f}{T_i}}\\
\Delta S=(2)(4200)\ln{\frac{333}{331}}=50.6\frac{J}{K}"